To analyze the alcohol content of a wine, a chemist needs 1.00 L of an aqueous 0.200 M K2Cr2O7 (potassium dichromate) solution. How much solid K2Cr2O7 must be weighted out to make this solution?


Describe also the laboratory procedures that the chemist must employ in preparing K2Cr2O7 solution.


1
Expert's answer
2022-03-29T07:48:03-0400

Сm(K2Cr2O7) = n(K2Cr2O7)/V(K2Cr2O7) = m(K2Cr2O7)/(Mr(K2Cr2O7)*V(K2Cr2O7))

m(K2Cr2O7) = Сm(K2Cr2O7)*Mr(K2Cr2O7)*V(K2Cr2O7)

Mr(K2Cr2O7) = 2*39 + 2*52 + 7*16 = 78 + 104 + 112 = 294 g/mol

m(K2Cr2O7) = 0.200 M * 294 g/mol * 1.00 L = 58.8 g


So, adding 58.8 g of potassium dichromate to enough water to make 1L will get a 0.2 M solution


sequence of actions:

  • weigh 58.8 g of K2Cr2O7 on electronic scales
  • pour K2Cr2O7 into a flask and add distilled water to the mark of 1 liter
  • stir with a glass rod
  • the solution is ready